OpenStudy (bebe):

what is 5/6 subtracted by 1/8

OpenStudy (amistre64):

what have you tried so far?

OpenStudy (aroub):

u have to find the lcd to subtract which is 48 ... 5/6-1/8 5/6*8 -1/8*6 40/48-6/48 34/48

OpenStudy (amistre64):

i find it rather simple to try to equate it to some unknown, call it N for number; clear the fractions, then solve for N

OpenStudy (amistre64):

\[\frac{5}{6}-\frac{1}{8}=N\text{ ; multiply by 6}\] \[5-\frac{6}{8}=6N\text{ ; reduce}\] \[5-\frac{3}{4}=6N\text{ ; multiply by 4}\] \[20-3=24N\text{ ; do the left side}\] \[17=24N\text{ ; divide off the 24}\] \[17/24 = N\]